The Company The best option decided to renew its security equipment valued at S/ 150,000 and financed it with the Bank BC
Answer #1

Total amount to be requested as loan to avail net S/150000 by ."best option " is ;

Let total amount requested as loan by bestoption is treated as 'X',

X = S/150000+ S/120+0.025X+0.1X+S/10000

X-0.025X-0.1X = S/160120

0.875X=S/160120

X=182994.286

Total amount to be requested by " bestoption" go avail a net bank loan of S/150000 is S/182995 (approx)

  

2 . Total cost borne by the bestoption to avail the loan is :

Calculation of equal semi annual payments is

Total amount requested by the best option is S/182,995

Interst rate is 20% per annum  

Total amount to be paid by the company after 18 months is

182995(1+20℅\2) ^3

=182995×1.331 = 243566

Let Equal semi annual installment is = P×r×(1+r) ^n÷(1+r) ^n -1

r= 20℅÷2 = 10% as it is semi annual payments

P= principal =182995

n=3 installment

182995×10%×(1+10%) ^3 ÷ (1+10%) ^3 -1

= 24356÷0.331=73585 equal installment semi annual

Total cost borne by the company due to lowering conditions is 120+ 10000+182995×2.5%+5489.85(note)

=20184.725

Note. Lost of interstbdue to with holding is 182995×10%×20%×18/12= 5489.85